# Semi-Supervised Preference Optimization with Limited Feedback

### v1

@@ D.2 Qualitative Comparison
In open-ended questions such as “What is the airspeed velocity of an unladen swallow?” (Table 11), SSPO stands out by maintaining a clear and concise structure while contextualizing the question accurately. It references the film “Monty Python and the Holy Grail,” distinguishes humor from fact, and provides an empirical estimate, resulting in an answer that is both informative and appropriately styled. In contrast, DPO and ORPO tend to over-elaborate with tangential details, while SimPO oversimplifies by dismissing the question altogether. In domain-specific prompts (from Table 12 to 15), SSPO continues to provide well-structured responses that surpass the superficial summaries of baselines, whether by presenting multi-step biological mechanisms with accurate terminology or by delivering comprehensive, actionable guides for brand marketing.
@@
DPO 1. Define your target audience: Know who your ideal customer is, what their needs and desires are, and how to reach them. 2. Create a strong brand identity: Develop a consistent brand image, message, and tone of voice that reflects your company’s values and resonates with your target audience. 3. Leverage social media: Use social media platforms like Facebook, Twitter, Instagram, and LinkedIn to connect with your audience, build brand awareness, and drive traffic to your website. 4. Use content marketing: Create valuable, informative, and engaging content like blog posts, videos, infographics, and podcasts to attract and retain customers. 5. Invest in search engine optimization (SEO): Optimize your website and content for search engines to improve your online visibility and attract more traffic. 6. Consider paid advertising: Use paid advertising like Google Ads, Facebook Ads, and Instagram Ads to reach a wider audience and drive more sales. 7. Measure and analyze your results: Use tools like Google Analytics to track your website traffic, social media engagement, and marketing campaigns to optimize your strategy and improve your ROI. 8. Stay consistent and authentic: Maintain a consistent brand image and message across all channels, and be authentic and transparent in your communication with customers.
ORPO 1. Define your target audience: Before you start promoting your brand, you need to know who your target audience is. You need to identify their age, gender, location, interests, and buying behavior. 2. Develop a strong brand identity: Your brand identity should be consistent across all platforms. Your logo, slogan, color scheme, and messaging should be recognizable and memorable. 3. Build a strong online presence: The internet is a powerful tool for brand promotion and marketing. You need to create a website, social media accounts, and other online platforms to reach your target audience. 4. Create valuable content: Content is king. You need to create valuable and relevant content that resonates with your target audience. This can include blog posts, videos, infographics, and other types of content. 5. Use social media effectively: Social media can be a powerful tool for brand promotion and marketing. You need to use social media platforms like Facebook, Instagram, Twitter, and LinkedIn to engage with your target audience and build brand awareness. 6. Leverage email marketing: Email marketing is an effective way to reach your target audience and promote your brand. You need to build an email list and send regular newsletters and promotions to your subscribers. 7. Partner with influencers: Influencer marketing can be a powerful way to reach a new audience and promote your brand. You need to find influencers who align with your brand values and work with them to create content and promotions. 8. Measure and track your results: It’s important to measure and track your results to determine the effectiveness of your brand promotion and marketing efforts. You can use tools like Google Analytics to track website traffic, social media engagement, and email open rates.
SimPO 1. Know your target audience: Understanding the needs, wants, and behavior of your target audience is crucial for effective brand promotion and marketing. 2. Create a strong brand identity: A strong brand identity, including a clear mission statement, memorable logo, and consistent messaging, helps to differentiate your brand from competitors and establish a strong presence in the market. 3. Develop a comprehensive marketing plan: A well-planned marketing strategy that includes a mix of traditional and digital marketing tactics, such as advertising, social media, email marketing, and content marketing, can help to reach and engage your target audience. 4. Leverage social media: Social media platforms, such as Facebook, Twitter, and Instagram, are powerful tools for brand promotion and marketing. Use these platforms to engage with your target audience, share valuable content, and build a community around your brand. 5. Provide exceptional customer service: Providing exceptional customer service not only helps to build loyalty and retention, but it also helps to generate positive word-of-mouth and online reviews, which can be powerful marketing tools. 6. Monitor and measure your results: Regularly monitoring and measuring the effectiveness of your marketing efforts is crucial for making data-driven decisions and continuously improving your brand promotion and marketing strategy.

### v2

@@ G.2 Qualitative Comparison
In open-ended questions such as “What is the airspeed velocity of an unladen swallow?” (Table 15), SSPO stands out by maintaining a clear and concise structure while contextualizing the question accurately. It references the film “Monty Python and the Holy Grail,” distinguishes humor from fact, and provides an empirical estimate, resulting in an answer that is both informative and appropriately styled. In contrast, DPO and ORPO tend to over-elaborate with tangential details, while SimPO oversimplifies by dismissing the question altogether. In domain-specific prompts (from Table 16 to 19), SSPO continues to provide well-structured responses that surpass the superficial summaries of baselines, whether by presenting multi-step biological mechanisms with accurate terminology or by delivering comprehensive, actionable guides for brand marketing.
@@
DPO 1. Defining the target audience: Understanding the demographics, needs, and preferences of the target market is crucial for creating effective brand promotions and marketing strategies. 2. Developing a strong brand identity: Establishing a consistent brand image, tone of voice, and message across all communication channels is essential for building brand awareness and recognition. 3. Utilizing various marketing channels: A mix of traditional and digital marketing channels such as social media, email marketing, content marketing, search engine optimization (SEO), pay-per-click (PPC) advertising, and public relations can help reach a wider audience and increase brand exposure. 4. Creating engaging content: Producing high-quality, relevant, and shareable content that resonates with the target audience can help generate buzz and increase brand awareness. 5. Building a community: Engaging with customers and building a community around the brand can help foster loyalty and increase brand advocacy. 6. Measuring and analyzing results: Regularly monitoring and analyzing the effectiveness of brand promotions and marketing strategies can help identify areas for improvement and optimize future campaigns.
ORPO 1. Define your target audience: Determine who your ideal customer is, their age, income, interests, and behavior patterns. This information will help you create effective marketing strategies that resonate with your target audience. 2. Create a unique brand identity: Your brand should stand out from the competition. Develop a distinct brand voice, logo, and visual identity that communicates your brand’s values and personality. 3. Develop a strong online presence: In today’s digital age, having a strong online presence is essential. Build a website, social media profiles, and other online channels that showcase your brand and engage with your target audience. 4. Build brand awareness: Use various marketing channels such as advertising, public relations, and content marketing to increase your brand’s visibility and reach. 5. Provide exceptional customer service: Providing excellent customer service is crucial for building brand loyalty and reputation. Ensure that your customers have positive experiences with your brand at every touchpoint. 6. Continuously measure and analyze your marketing efforts: Use analytics tools to measure the effectiveness of your marketing campaigns and make data-driven decisions to optimize your strategies. 7. Stay true to your brand values: Your brand’s values should be the foundation of all your marketing and promotional efforts. Ensure that your messaging and marketing strategies align with your brand’s mission and purpose.
SimPO 1. Know Your Target Audience: Determine your ideal customer and what drives their purchasing decisions. 2. Develop a Strong Brand Identity: Create a clear and consistent brand message that resonates with your audience. 3. Build a Strong Online Presence: Utilize social media, a website, and other online channels to reach your audience and establish your brand. 4. Provide Exceptional Customer Service: Provide excellent customer service to build loyalty and attract repeat business. 5. Leverage Influencer Marketing: Partner with influencers in your industry to reach a wider audience and increase brand awareness. 6. Monitor and Measure Results: Track and analyze your marketing efforts to understand what works and what doesn’t, and adjust your strategy accordingly. 7. Offer Value: Provide value to your customers through promotions, discounts, and other incentives to encourage them to engage with your brand. 8. Collaborate with Other Brands: Collaborate with complementary brands to reach new audiences and increase brand visibility. 9. Stay Adaptable: Stay up-to-date on industry trends and adapt your marketing strategy to remain relevant and effective.